OpenStudy (anonymous):

Given the quation:4x+3y=-12 write the euation in slope-intercept form.

OpenStudy (anonymous):

Slope-intercept form is y = mx + b. So solve for y. 4x + 3y = -12 -4x 3y = -4x - 12 /3 y = -4/3x - 4

OpenStudy (anonymous):

thank you sooo much what about writing that same equation in function notation?